OpenStudy (anonymous):

Find the solution of y = –2x – 6 for x = 2?

OpenStudy (anonymous):

y = -2(2) - 6 y = -4 - 6 y = -10 sub y = -10 back into the equation to find x -10 = -2x - 6 2x =10 - 6 2x = 4 x = 2 solution is x = 2 and y = -10 so (2, -10)

OpenStudy (anonymous):

y= -2(2) -6 _2(2)= 2 x -2 = -4 -4 -6 = -10 y=-10